# Higher-dimensional flying wing Steady Ricci Solitons

### v1

@@ 4.1. Relation between Ricci flow and Ricci-DeTurck flow
see the computation in [41, Lemma 2.1] for example. Here we use the convention $\mathrm{sec}_{ij}=R_{ijij}$ in orthogonal frame.
@@
where $W^{k}=g^{ij}\left(\Gamma_{ij}^{k}-\tilde{\Gamma}_{ij}^{k}\right)$ is a time dependent vector field, then $\hat{g}(t):=\Psi_{t}^{*}g(t)$ is also a solution to the Ricci flow with $\hat{g}(T)=g(T)$. If $g(t)$ is smooth at $t=0$, we might choose $T=0$.
@@ 4.3. Evolution equation
where we have used $\mathrm{Rm}(\tilde{g})\geq-1$ and [7, Lemma 3.2].
@@ 4.4. Some a-priori estimates
    $\mathrm{Rm}(\tilde{g}(t))\geq-r^{-2}$ on $M\times[0,T]$.
@@
We might assume $r=1$ by scaling. We follow [44, Lemma 7.1] to construct cut-off function. Let $\phi$ be a smooth non-increasing function on $0,+\infty)$ such that $\phi\equiv 1$ on $0,\frac{3}{2}]$, vanishes outside $[0,2]$ and satisfies $|\phi^{\prime}|^{2}\leq 10^{3}\phi,\phi^{\prime\prime}\geq-10^{3}\phi$. Define $\varphi(x,t):=e^{-10^{3}t}\phi\left({d_{\tilde{g}(t)}(x_{0},x)+C_{n}\sqrt{{\alpha}t}}\right)$ so that
@@
in the sense of barrier, and thus distribution sense [38, Appendix]. Considering
@@
    $\mathrm{Rm}(\tilde{g}(t))\geq-r^{-2}$ on $M\times[0,T]$.
@@
This follows from a straightforward modification of [34, Lemma 2.6] using (4.8), with extra attention to the smallness of $g-\hat{g}$. ∎
@@ 4.5. Heat kernel estimates
Suppose $\tilde{g}(t)$ is a complete Ricci flow on $M\times(0,T]$ such that (a) and (b) above hold, then there is $C_{1}(n,{\alpha})>0$ such that the heat kernel $G(x,t;y,s)$ with respect to operator $\partial_{t}-\Delta_{\tilde{g}(t)}-R_{\tilde{g}(t)}$ satisfies
@@
We now choose $r$ suitably to draw conclusion when $d_{\tilde{g}(0)}(x,y)\geq 2C_{3}\sqrt{t_{0}}$ for large $C_{3}(n,{\alpha})>0$. By distance distortion [43, Lemma 3.1], $d_{\tilde{g}(s)}(x,y)\geq C_{3}\sqrt{t_{0}}$ for $s\in[0,t_{0}/2]$. In this case, we choose $r=10^{-3}C_{3}\sqrt{t_{0}}$. If $(z,\tau)\in\partial\Omega\times[0,t_{0}/2]$, then
@@ 4.6. Proof of weak stability
on $\Omega\times(0,T]$, then for $(x_{0},t_{0})\in\Omega\times(0,T\wedge 1]$,
@@
The case when $(M,\tilde{g}(t))$ is complete non-compact with bounded curvature up to $t=0$ is similar to the compact case. We focus on the case when $\tilde{g}(t)$ has no uniform bounded curvature up to $t=0$. We let $\phi_{i}$ be a sequence of smooth cut-off function supported on $B_{\tilde{g}(0)}(x_{0},i)$ which exhausts $M$. Let $\sigma_{i}\to 0$ be a sequence of time and we approximate $g_{0}$ by $g_{i,0}:=\phi_{i}g_{0}+(1-\phi_{i})\tilde{g}(\sigma_{i})$ which has bounded curvature. We choose $\sigma_{i}$ so that $|g_{i,0}-\tilde{g}(\sigma_{i})|_{\tilde{g}(\sigma_{i})}\leq 2\varepsilon$. By the work of Simon [42] (see also the work of Shi [41]), there exists a short-time solution $g_{i}(t)$ to (4.3) with respect to $\tilde{g}(\sigma_{i}+t)$ starting from $g_{i,0}$. Using the same analysis as in the compact case, we see that $g_{i}(t)$ exists up to $T\wedge 1$ and satisfies $|g_{i}(t)-\tilde{g}(\sigma_{i}+t)|\leq 4C_{1}\varepsilon$, provided that $\bar{\varepsilon}_{0}\leq(20C_{1}C_{10})^{-1}$. This also enables us to let $i\to+\infty$ to obtain the global solution $g(t)$ by interior estimates of Ricci-DeTurck flow, see [42, Lemma 2.4] or [14, Proposition 2.2] for example. This proves the quantitative existence in all cases. The $C^{0}_{loc}$ convergence as $t\to 0$ follows from the proof [42, Theorem 5.2] where the argument is purely local.
@@
By [43, Lemma 3.1], the Hausdorff measure $\mathcal{H}^{n-1}_{\tilde{d}_{0}}$ of $\tilde{d}_{0}$ satisfies $\mathcal{H}_{d_{\tilde{g}(t)}}\leq\mathcal{H}_{\tilde{d}_{0}}$ for all $t\in(0,S]$. Using this, the heat kernel estimate from Proposition 4.8 and the assumed estimate of $h$,
@@
By [43, Lemma 3.1], the distance function $\tilde{d}_{0}$ generates the same topology as $M$.
@@ 5.1. Ricci flow smoothing of $\mathcal{S}_{j}$ by stability
There exists $\delta_{0}>0$ depending only on $j$ such that the following holds: Suppose for $\hat{\beta}\in[j^{-1},1]^{n-1}$ and $\tau>0$, we have already constructed a Ricci flow $g_{\hat{\beta},\tau}(t)$ coming out of $g_{0,\hat{\beta},\tau}$ and there exist a homeomorphism $G_{0,\hat{\beta},\tau}$ and a finite set $\mathtt{S}_{\hat{\beta},\tau}$ such that
@@
for $t\in(0,S]$. By Lemma 4.7, we also know $\Phi_{{\beta},\hat{\beta},\tau}(t)$ converges to some continuous map $\Phi_{0,{\beta},\hat{\beta},\tau}$ as $t\to 0$. By interior estimate, see [14, Proposition 2.2] for example, $\Phi_{0,{\beta},\hat{\beta},\tau}$ is smooth outside $\Phi_{0,{\beta},\hat{\beta},\tau}^{-1}\left(\mathtt{S}_{\hat{\beta},\tau}\cup G_{0,\hat{\beta},\tau}^{-1}(\mathtt{Tips})\right)$ where the convergence as $t\to 0$ is in local smooth topology. Consequently, $g_{{\beta},\hat{\beta},\tau}(t)$ converges in $C^{\infty}_{\text{loc}}$ to $(\Phi_{0,{\beta},\hat{\beta},\tau})^{*}(G_{0,\hat{\beta},\tau})^{*}g_{0,{\beta},\hat{\beta},\tau}$ as $t\to 0$ away from $\Phi_{0,{\beta},\hat{\beta},\tau}^{-1}\left(\mathtt{S}_{\hat{\beta},\tau}\cup G_{0,\hat{\beta},\tau}^{-1}(\mathtt{Tips})\right)$. Furthermore we have
@@
Since we have shown $\text{\rm Rm}(g_{{\beta},\hat{\beta},\tau}(t))\geq 1$, by the same argument as in the proof of (vii) in Theorem 3.2, we see that the Ricci flow $g_{{\beta},\hat{\beta},\tau}(t)$ satisfies (i) and (ii), for the same ${\alpha}>0$ depending only on $\mathrm{Vol}(\mathbb{S}^{n-1},g_{0,(j^{-1},...,j^{-1})})$. Furthermore, it follows from [43, Lemma 3.1] and
@@
We denote distance $d_{0,{\alpha}}:=\lim_{t\to 0^{+}}d_{g_{\alpha}(t)}$ and $d_{0,k}:=\lim_{t\to 0^{+}}d_{g_{k}(t)}$ which exist by [43, Lemma 3.1]. We next claim that $\mathtt{S}_{\alpha}$ is of zero (n1)(n-1)-Hausdorff measure with respect to distance $d_{0,{\alpha}}:=\lim_{t\to 0^{+}}d_{g_{\alpha}(t)}$. By construction, $\Phi_{0,{\alpha},0}$ is a $(1+\bar{\varepsilon}_{0})^{1/2}$ bi-Lipschitz map from $(\mathbb{S}^{n-1},d_{0,{\alpha}})$ to $(\mathbb{S}^{n-1},d_{0,N_{\alpha}})$. Hence,
@@
Claim 5.2.
For any $\varepsilon>0$, there exists $\delta>0$ so that if $|{\alpha}-{\beta}|<\delta$ and ${\alpha},{\beta}\in[j^{-1},1]^{n-1}$, then there exists a solution to the Ricci-harmonic map heat flow
(5.27) $\left\{\begin{array}[]{ll}\partial_{t}\Psi_{{\alpha},{\beta}}=\Delta_{g_{\beta}(t),g_{\alpha}(t)}\Psi_{{\alpha},{\beta}},\;\;\text{for}\;t\in(0,S];\\[2.84526pt] \lim_{t\to 0}\Psi_{{\alpha},{\beta}}(t):=\psi_{0,{\alpha}}^{-1}\circ\psi_{0,{\beta}}\end{array}\right$.
such that
on $(0,S]$.
@@
for all $t\in[t_{i},S\wedge T_{max}]$. This implies high order regularity of $\Psi_{{\alpha},{\beta},i}$ by [34, Proposition 2.1] and $T_{max}=S$. In particular, we might pass $i\to+\infty$ and obtain a solution $\Psi_{{\alpha},{\beta}}$ to (5.27). The convergence as $t\to 0$ follows from a similar argument in (5.12):
@@
for $0<s<t$. By letting $s\to 0$ using [43, Lemma 3.1], we show the convergence to the initial data as $t\to 0$. Moreover, the convergence is smooth outside $\mathtt{S}_{\beta}\cup\Psi_{{\alpha},{\beta}}(0)^{-1}(\mathtt{S}_{\alpha})$ by interior estimate of Ricci-DeTurck flow. This proves the existence of solution to (5.27) for the case of $k=0$.
@@
The lower bound is similar so that

### v2

@@ 4.1. Relation between Ricci flow and Ricci-DeTurck flow
see the computation in [43, Lemma 2.1] for example. Here we use the convention $\mathrm{sec}_{ij}=R_{ijij}$ in orthogonal frame.
@@
where $W^{k}=g^{ij}\left(\Gamma_{ij}^{k}-\tilde{\Gamma}_{ij}^{k}\right)$ is a time dependent vector field, then $\hat{g}(t):=\Psi_{t}^{*}g(t)$ is also a solution to the Ricci flow with $\hat{g}(T)=g(T)$. If $g(t)$ is smooth at $t=0$, we might choose $\Psi_{0}=\mathrm{Id}$.
@@ 4.3. Evolution equation
where we have used $\mathrm{sec}(\tilde{g})\geq-1$. Here we have used the fact that $2\text{\rm Ric}(R)\leq\mathrm{scal}(R)$ for curvature tensor $R$ with $\mathrm{sec}(R)\geq 0$, see the proof of [9, Lemma 3.2].
@@ 4.4. Some a-priori estimates
    $\mathrm{sec}(\tilde{g}(t))\geq-r^{-2}$ on $M\times[0,T]$.
@@
We might assume $r=1$ by scaling. We follow [46, Lemma 7.1] to construct cut-off function. Let $\phi$ be a smooth non-increasing function on $0,+\infty)$ such that $\phi\equiv 1$ on $0,\frac{3}{2}]$, vanishes outside $[0,2]$ and satisfies $|\phi^{\prime}|^{2}\leq 10^{3}\phi,\phi^{\prime\prime}\geq-10^{3}\phi$. Define $\varphi(x,t):=e^{-10^{3}t}\phi\left({d_{\tilde{g}(t)}(x_{0},x)+C_{n}\sqrt{{\alpha}t}}\right)$ so that
@@
in the sense of barrier, and thus distribution sense [40, Appendix]. Considering
@@
    $\mathrm{sec}(\tilde{g}(t))\geq-r^{-2}$ on $M\times[0,T]$.
@@
This follows from a straightforward modification of [36, Lemma 2.6] using (4.8), with extra attention to the smallness of $g-\hat{g}$. ∎
@@ 4.5. Heat kernel estimates
Suppose $\tilde{g}(t)$ is a complete Ricci flow on $M\times(0,T]$ such that
  1. (a)
    $|\text{\rm Rm}(\tilde{g}(t))|\leq{\alpha}t^{-1}$;
  2. (b)
    $\mathrm{inj}(\tilde{g}(t))\geq\sqrt{{\alpha}^{-1}t}$
for some ${\alpha}>0$, then there is $C_{1}(n,{\alpha})>0$ such that the heat kernel $G(x,t;y,s)$ with respect to operator $\partial_{t}-\Delta_{\tilde{g}(t)}-R_{\tilde{g}(t)}$ satisfies
@@
We now choose $r$ suitably to draw conclusion when $d_{\tilde{g}(0)}(x,y)\geq 2C_{3}\sqrt{t_{0}}$ for large $C_{3}(n,{\alpha})>0$. By distance distortion [45, Lemma 3.1], $d_{\tilde{g}(s)}(x,y)\geq C_{3}\sqrt{t_{0}}$ for $s\in[0,t_{0}/2]$. In this case, we choose $r=10^{-3}C_{3}\sqrt{t_{0}}$. If $(z,\tau)\in\partial\Omega\times[0,t_{0}/2]$, then
@@ 4.6. Proof of weak stability
and $\mathrm{sec}(\tilde{g}(t))\geq-1$ on $\Omega\times(0,T]$, then for $(x_{0},t_{0})\in\Omega\times(0,T\wedge 1]$,
@@
The case when $(M,\tilde{g}(t))$ is complete non-compact with bounded curvature up to $t=0$ is similar to the compact case. We focus on the case when $\tilde{g}(t)$ has no uniform bounded curvature up to $t=0$. We let $\phi_{i}$ be a sequence of smooth cut-off function supported on $B_{\tilde{g}(0)}(x_{0},i)$ which exhausts $M$. Let $\sigma_{i}\to 0$ be a sequence of time and we approximate $g_{0}$ by $g_{i,0}:=\phi_{i}g_{0}+(1-\phi_{i})\tilde{g}(\sigma_{i})$ which has bounded curvature. We choose $\sigma_{i}$ so that $|g_{i,0}-\tilde{g}(\sigma_{i})|_{\tilde{g}(\sigma_{i})}\leq 2\varepsilon$. By the work of Simon [44] (see also the work of Shi [43]), there exists a short-time solution $g_{i}(t)$ to (4.3) with respect to $\tilde{g}(\sigma_{i}+t)$ starting from $g_{i,0}$. Using the same analysis as in the compact case, we see that $g_{i}(t)$ exists up to $T\wedge 1$ and satisfies $|g_{i}(t)-\tilde{g}(\sigma_{i}+t)|\leq 4C_{1}\varepsilon$, provided that $\bar{\varepsilon}_{0}\leq(20C_{1}C_{10})^{-1}$. This also enables us to let $i\to+\infty$ to obtain the global solution $g(t)$ by interior estimates of Ricci-DeTurck flow, see [44, Lemma 2.4] or [16, Proposition 2.2] for example. This proves the quantitative existence in all cases. The $C^{0}_{loc}$ convergence as $t\to 0$ follows from the proof [44, Theorem 5.2] where the argument is purely local.
@@
By [45, Lemma 3.1], the Hausdorff measure $\mathcal{H}^{n-1}_{\tilde{d}_{0}}$ of $\tilde{d}_{0}$ satisfies $\mathcal{H}_{d_{\tilde{g}(t)}}\leq\mathcal{H}_{\tilde{d}_{0}}$ for all $t\in(0,S]$. Using this, the heat kernel estimate from Proposition 4.8 and the assumed estimate of $h$,
@@
By [45, Lemma 3.1], the distance function $\tilde{d}_{0}$ generates the same topology as $M$.
@@ 5.1. Ricci flow smoothing of $\mathcal{S}_{j}$ by stability
There exists $\delta_{0}>0$ depending only on $j$ such that the following holds: Suppose for $\hat{\beta}\in[j^{-1},1]^{n-1}$ and $\tau>0$, we have already constructed a Ricci flow $g_{\hat{\beta},\tau}(t),t\in(0,S]$ coming out of $g_{0,\hat{\beta},\tau}$ and there exist a homeomorphism $G_{0,\hat{\beta},\tau}$ and a finite set $\mathtt{S}_{\hat{\beta},\tau}$ such that
@@
for $t\in(0,S]$. By Lemma 4.7, we also know $\Phi_{{\beta},\hat{\beta},\tau}(t)$ converges to some continuous map $\Phi_{0,{\beta},\hat{\beta},\tau}$ as $t\to 0$. By interior estimate, see [16, Proposition 2.2] for example, $\Phi_{0,{\beta},\hat{\beta},\tau}$ is smooth outside $\Phi_{0,{\beta},\hat{\beta},\tau}^{-1}\left(\mathtt{S}_{\hat{\beta},\tau}\cup G_{0,\hat{\beta},\tau}^{-1}(\mathtt{Tips})\right)$ where the convergence as $t\to 0$ is in local smooth topology. Consequently, $g_{{\beta},\hat{\beta},\tau}(t)$ converges in $C^{\infty}_{\text{loc}}$ to $(\Phi_{0,{\beta},\hat{\beta},\tau})^{*}(G_{0,\hat{\beta},\tau})^{*}g_{0,{\beta},\hat{\beta},\tau}$ as $t\to 0$ away from $\Phi_{0,{\beta},\hat{\beta},\tau}^{-1}\left(\mathtt{S}_{\hat{\beta},\tau}\cup G_{0,\hat{\beta},\tau}^{-1}(\mathtt{Tips})\right)$. Furthermore we have
@@
Since we have shown $\text{\rm Rm}(g_{{\beta},\hat{\beta},\tau}(t))\geq 1$, by the same argument as in the proof of (vii) in Theorem 3.2, we see that the Ricci flow $g_{{\beta},\hat{\beta},\tau}(t)$ satisfies (i) and (ii), for the same ${\alpha}>0$ depending only on $\mathrm{Vol}(\mathbb{S}^{n-1},g_{0,(j^{-1},...,j^{-1})})$. Furthermore, it follows from [45, Lemma 3.1] and
@@
We denote distance $d_{0,{\alpha}}:=\lim_{t\to 0^{+}}d_{g_{\alpha}(t)}$ and $d_{0,k}:=\lim_{t\to 0^{+}}d_{g_{k}(t)}$ which exist by [45, Lemma 3.1]. We next claim that $\mathtt{S}_{\alpha}$ is of zero (n1)(n-1)-Hausdorff measure with respect to distance $d_{0,{\alpha}}:=\lim_{t\to 0^{+}}d_{g_{\alpha}(t)}$. By construction, $\Phi_{0,{\alpha},0}$ is a $(1+\bar{\varepsilon}_{0})^{1/2}$ bi-Lipschitz map from $(\mathbb{S}^{n-1},d_{0,{\alpha}})$ to $(\mathbb{S}^{n-1},d_{0,N_{\alpha}})$. Hence,
@@
for all $t\in[t_{i},S\wedge T_{max}]$. This implies high order regularity of $\Psi_{{\alpha},{\beta},i}$ by [36, Proposition 2.1] and $T_{max}\geq S$. In particular, we might pass $i\to+\infty$ and obtain a solution $\Psi_{{\alpha},{\beta}}$ to (5.27). The convergence as $t\to 0$ follows from a similar argument in (5.12):
@@
for $0<s<t$. By letting $s\to 0$ using [45, Lemma 3.1], we show the convergence to the initial data as $t\to 0$. Moreover, the convergence is smooth outside $\mathtt{S}_{\beta}\cup\Psi_{{\alpha},{\beta}}(0)^{-1}(\mathtt{S}_{\alpha})$ by interior estimate of Ricci-DeTurck flow. This proves the existence of solution to (5.27) for the case of $k=0$.
@@
The lower bound is similar, so that
@@
We end this subsection by presenting another application of our stability result. Using the method in the proof of Theorem 5.2, we are able to show that the expander constructed by Deruelle [19] is stable with respect to $L^{\infty}$ variation of the links. Quantitatively, we have the following:
Theorem 5.4.
For any $n\in\mathbb{N}$ and $v>0$, there exists $\delta_{0}(n,v),\Lambda(n,v)>0$ such that the following holds. Suppose $X_{i}:=(\mathbb{S}^{n-1},h_{i})$ where $i=1,2$ are two smooth manifolds such that
  1. (i)
    $\mathrm{Vol}_{h_{1}}(X_{1})\geq v>0$;
  2. (ii)
    $\mathrm{Rm}(h_{i})>1$ for $i=1,2$;
  3. (iii)
    $(1-\varepsilon)h_{2}\leq h_{1}\leq(1+\varepsilon)h_{2}$ for some $0<\varepsilon<\delta_{0}$.
Let $(M_{i},g_{i},f_{i},p_{i})$ be the unique expanding Ricci solitons coming out of $C(X_{i})$, constructed by Deruelle [19] for $i=1,2$, then there exists a diffeomorphism $\Phi:M_{1}\to M_{2}$ such that
(5.42) $(1-\Lambda\varepsilon)g_{1}\leq\Phi^{*}g_{2}\leq(1+\Lambda\varepsilon)g_{1},\;\;\text{on}\;M_{1}.$
Furthermore for any $k\in\mathbb{N}$, there exists $C_{k}(n,v)>0$ so that
$\sup_{M_{1}}|\nabla^{k,g_{1}}(\Phi^{*}g_{2}-g_{1})|\leq C_{k}\varepsilon.$
Proof.
Since the proof follows from modification of that in claim 5.1, we only give a sketch. It suffices to show the quantitative stability. To do this, we construct $(M_{2},g_{2},f_{2},p_{2})$ using an alternative way.
We let $g_{1}(t)$ be the Ricci flow associated to $(M_{1},g_{1},f_{1},p_{1})$. By identifying the distance limit of $d_{g_{1}(t)}$ as $t\to 0$ with $C(X_{1})$ using [45, Lemma 3.1], we might assume $M_{1}\equiv C(X_{1})$ where $g_{2,0}:=dr^{2}+r^{2}h_{2}$ is also a $L^{\infty}$ metric on $M_{1}$ such that
(5.43) $(1-\varepsilon)g_{2,0}\leq dr^{2}+r^{2}h_{1}\leq(1+\varepsilon)g_{2,0}$
on $M_{1}$. We follow the method in the proof of claim 5.1. For $\sigma\to 0^{+}$, we consider the $L^{\infty}$ metric
(5.44) $g_{2,0,\sigma}:=\phi_{\sigma}g_{1}(\sigma)+(1-\phi_{\sigma})g_{2,0}$
where $\phi_{\sigma}(x):=\phi(d_{g_{1}(\sigma)}(x,o_{tips})/\sigma^{1/4})$.
As $\sigma\to 0$, $g_{2,0,\sigma}\to g_{2,0}$ smoothly outside the tips. Arguing in a similar way as the construction of $\hat{g}_{{\beta},\hat{\beta},\tau,\sigma}$ in the proof of claim 5.1, this will enable us to apply Theorem 4.1 (with rescaling) to $g_{2,0,\sigma}$ and obtain a Ricci-DeTurck flow $\hat{g}_{2,\sigma}(t)$ with respect to $g_{1}(\sigma+t)$ starting from $g_{2,0,\sigma}$ such that
(5.45) $\left(1-\Lambda\varepsilon\right)g_{1}(\sigma+t)\leq\hat{g}_{2,\sigma}(t)\leq\left(1+\Lambda\varepsilon\right)g_{1}(\sigma+t)$
on $M_{1}\times(0,+\infty)$, for all $\sigma\to 0^{+}$. By passing $\sigma\to 0$ subsequently and Lemma 4.7, we obtain a smooth solution to the Ricci-DeTurck flow $\hat{g}_{2}(t)$ with respect to $g_{1}(t)$ on $M_{1}\times(0,+\infty)$ satisfying
(5.46) $\left(1-\Lambda\varepsilon\right)g_{1}(t)\leq\hat{g}_{2}(t)\leq\left(1+\Lambda\varepsilon\right)g_{1}(t).$
Moreover by [16, Proposition 2.2], $\hat{g}_{2}(t)$ is smooth up to $t=0$ with $\hat{g}_{2}(0)=g_{2,0}$ outside the tips.
Now we might use the argument in the proof of claim 5.1 to show that indeed $0<\mathrm{Rm}(\hat{g}_{2}(t))\leq{\alpha}t^{-1}$ for some ${\alpha}(n,v)>0$, using crucially the fact that $\delta_{0}$ is small and the singularity is isolated. By considering the Ricci-DeTurck ODE as in (5.10), we obtain a Ricci flow $\tilde{g}_{2}(t)$ coming out of $C(X_{2})$ in the pointed Gromov-Hausdorff sense, thanks to [45, Lemma 3.1]. By [10, Theorem 6.1], $\tilde{g}_{2}(1)$ is a gradient expanding Ricci soliton coming out of $C(X_{2})$. Since the link $X_{2}$ is smooth, by [19, Theorem 1.3], the new solution $(M_{1},\tilde{g}_{2}(1))$ is isometric to the expander $(M_{2},g_{2})$ constructed by Deruelle. The claimed estimates follow from (5.46) and Lemma 4.7. ∎